What is the present value of $1,600 per year, at a discount rate of 7 percent, if the first payment will be received 7 years from now and the last payment will be received 30 years from now? Hint: draw a timeline and count the number of payments.
1
Expert's answer
2012-10-02T07:30:10-0400
1600*1.07^7 =2 569$ will be in 7 years.
1600*1.07^30=12 179 $ will be in 30 years
if the first payment is received 7 years from now then in 30 years it'll be
1600*1.07^(30-7)= 7 585 $